300 298 307 279 344 ?
Answer by Edwin McCravy(20056)

Start with 1st term 300.

Subtract 1 more than 13, which is 1+1=2, get 300-2 = 298.
     Add 1 more than 23, which is 8+1=9, get 298+9 = 307.
Subtract 1 more than 33, which is 27+1=28, get 307-28 = 279.
     Add 1 more than 43, which is 64+1=65, get 279+65 = 344.
Subtract 1 more than 53, which is 125+1=126, get 344-126 = 218*.
     Add 1 more than 63, which is 216+1=217, get 218+217 = 435.

"Answer = 218

Edwin